OpenStudy (anonymous):

-x^2+10x-16 find the zeros of the polynomial

OpenStudy (anonymous):

Are you studying the quadratic equation?

OpenStudy (zzr0ck3r):

devide away the - x^2 -10x + 16 = 0 (x-5)^2 = 9 x-5 = +-sqrt(9 x = 5 +- sqrt(9) x = 5 + sqrt(9) and x = 5 - sqrt(9) this is the method of completing the square

OpenStudy (anonymous):

x = 5 + sqrt(9) =8 and x = 5 - sqrt(9)=2

OpenStudy (campbell_st):

why not just factorise -(x -2)(x -8) = 0 then x = 2, 8

OpenStudy (zzr0ck3r):

I complete the square everytime, because it always works and is super easy.